10. I bought 1,000 shares of stock for $5 ea two years ago. I sold them this year for $10 ea. I also bought 500 shares of a different stock 3 years ago paying $9 ea. I sold them this year for $8 ea. My income is $100,000 per year (I am married). How much capital gains tax do I owe, if any? (Answer is a dollar amount or a zero).
Answer #1

LTCG on 1,000 shares of stock= (10-5)*1000= $5,000

LTCL on 500 shares of stock= (9-8)*500= $500

LTCL can be set off against LTCG, thus taxable LTCG= 5,000-500= $4,500

For income in the range of $38,601 to $425,800, tax is applicable @15%,

thus tax on LTCG= 4,500*15%= $675